A 197 g cube of ice at 0◦C is dropped into 1.4 kg of water that was originally at 78◦C. What is the final temperature of the water after after the ice melts? The latent heat of fusion of water is 3.33 × 105 J/kg. Answer in units of ◦C.
Answer #1

mass of ice = 197 g

mass of water = 1.4 kg

initial temperature is 78 C

let final temperature be T C

initial temperature of ice = 0 C

heat lost by water = heat gained by ice

MC dt=mL+mCdt'

C is specific heat of water = 1 ca;/g

L= latent heat of ice = 80 cal/g

dt=78.-T

dt'=T-0=T

1400 x 1 x (78-T)=(197 x 80) + (197 x 1 x T)

109200-1400 T=15760+197 T

T= 58.509 C